Julia, who is 1.51 m tall, wishes to find the height of a tree with a shadow 31.05 m long. She walks 20.70 m from the base of the tree along the shadow of the tree until her head is in a position where the tip of her shadow exactly overlaps the end of the tree top's shadow. How tall is the tree? Round to the nearest hundredth.
Answer by checkley77(12844)

1.51/x=(31.05-20.7)/31.05
1.51/x=10.35/31.05
10.35x=1.51*31.05
10.35x=46.8855
x=46.8855/10.35
x=4.53 m. for the height of the tree.
P500f:
1.51/4.53=10.35/31.05
1/3=1/3